OpenStudy (anonymous):

Find a and b such that the derivative of f exists everywhere, if f(x)= {x^2+3, (x < or =1) and ax^3+bx+4, x>1

OpenStudy (anonymous):

this is much easier than it looks take the derivative of both replace \(x\) by \(1\) and set them equal

OpenStudy (anonymous):

so 2x and a3x^2+b 2(1)=1 a3(1)^2+b=1 ?

OpenStudy (anonymous):

lost.

OpenStudy (anonymous):

still trying to figure out if I am on the right track. I am lost right now. No idea what to do after I find their derivatives never had an a and b in the equation like this.

OpenStudy (anonymous):

@Fakshon as far as derivatives are concerned...you got that part right.. 2x and 3a*x^2+b now you have to set: 3a*x^2+b=2x and put x=1

OpenStudy (anonymous):

so you got one equation with a and b?

OpenStudy (anonymous):

still lost?Just nod...we can go to the begining

OpenStudy (anonymous):

you function is defined in 2 different ways depending on the interval. The limit point of the intervals is x=1. Independently, the derivative exist on each interval separatly, but for them to exist in all R it is needed that they are equal in this limiting point x=1. So like @satellite73 said, find first each derivative. Later set x=1 and put them equal. What is left is to choose a and b that make the equality happen. That's it

OpenStudy (anonymous):

@myko we get one equation in a and b by differentiating both parts and setting them equal at x=1...the other equation is obtained from the fact that function is continuous at x=1...since continuous functions are differentiable and vice versa... so we can equalize x^2+3 and ax^3+bx+4 at x=1..so that the function is continuous.. now two equations..two unknowns.. can be worked out

OpenStudy (anonymous):

right